OpenStudy (anonymous):

5/13+5/26 in a simplified fraction or whole number

OpenStudy (anonymous):

15/26

OpenStudy (anonymous):

15/26

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

2/2(5/13)= 10/26+5/26= 15/26

OpenStudy (jamesj):

C'mon. Use the method we just developed in the last problem to figure this out. If you put all your homework problems on here and just copy the answers, you will learn zero and fail the exam. I promise. So c'mon.

OpenStudy (anonymous):

your right I should at least try

jimthompson5910 (jim_thompson5910):

they need that as a disclaimer all over this site lol

OpenStudy (jamesj):

What you could do is try the next one by yourself and then ask us if the answer you arrived it is correct.

OpenStudy (anonymous):

is that the LCM of 5 and 13

OpenStudy (anonymous):

im glad u said that i finally am able to understand it on my own I tried this one 1/6+1/7= 7/42+6/42=13/42